# Yuri Balashov

### v1

@@ 7.1 Selective Ablation of Fine-Tuning Data
One direct way to probe the origins of LLMs’ translation behavior is to selectively remove or add certain data from an LLM’s fine-tuning corpus and examine the impact. As outlined above (Section 6.2), this “retrospective surgery” approach was pioneered by Briakou et al. (2023) on PaLM’s smaller replicas in a full-scale training regime. But it could be done in a less expensive fine-tuning mode and applied to other, increasingly available open-source models. For example, one could take an existing pre-trained model, choose a new language pair, and continue training or fine-tuning the model on an additional dataset from which all parallel sentences for a new language pair are ablated, and then do the same with a version of that dataset where those pairs are exclusively retained but all the monolingual portions in the two chosen languages are removed. By comparing these two extremes—“no parallel data” vs. “only parallel data”—one could, in principle, observe stark differences in resulting translation performance. The dualism hypothesis predicts that a model fine-tuned without any parallel examples would still develop some translation ability in a new language pair (via global semantic alignment), but it might translate more slowly or less idiomatically for complex sentences. Conversely, a model fine-tuned only on parallel sentences would certainly learn direct mappings but might struggle on inputs that require world knowledge or cross-sentence context (since the training signal was confined to one-to-one sentence translations). If adding parallel data for, say, Russian-Japanese suddenly causes the model to produce very literal, consistent Japanese translations for Russian inputs (while translation outputs for other relevant language pairs remain unchanged), that would be a strong indication of a local learning effect.
Such controlled fine-tuning experiments could be done on smaller proxy models first, and data could be added incrementally, in both regimes. This might be similar to classic studies in cross-lingual word embeddings where parallel data were incrementally added to chiefly monolingual training to observe alignment effects (e.g. refining mBERT’s semantic space; cf. Conneau et al. 2020b). Key data sources for real-world ablation studies include multilingual web crawls like mC4 or Common Crawl derivatives, where one can identify and remove bilingual segments (using language detectors and alignment algorithms). However, smaller corpus resources—for example, from the latest MT benchmarks or recent real-life translation projects (e.g., the Reeve Trilingual Corpus, Balashov et al. 2025)—could also be used in the fine-tuning mode. Model sizes should vary: experiments on 1B, 5B, and 10B parameter models can indicate scaling trends.
@@ 7.2 Synthetic Tasks and Controlled Probes
Another strategy is to design diagnostic tasks that isolate local vs. global translation skills without needing to retrain or even fine-tune the model. One intriguing idea is to present the model with synthetic bilingual contexts at inference time to see how it leverages them. For instance, we could feed the model a prompt that includes a few embedded parallel examples for a new language pair (e.g. a mini-glossary or several translated sentences) and then ask it to translate a novel sentence. This is akin to few-shot prompting. If local learning is an operative mechanism, the model should significantly improve or change style given those explicit examples—essentially performing cross-lingual “translation memory” look-up—whereas if it relies mostly on global knowledge, a few extra examples might not change its performance much. By varying the content of these examples (literal translations vs. paraphrastic ones) and observing the output, one can infer the model’s default bias. A model heavily skewed to local learning might even mistranslate by over-following the format of a misleading example (e.g., if given a bad literal translation as a cue, it might mimic that literality). In contrast, a globally oriented model would stick closer to conveying meaning and be less perturbed by small prompt changes. Recent work on prompt-based analysis of LLM translation indicates that certain prompt formats act as “soft switches” for translation style (Aldosari and Altuwairesh, 2025); this can be leveraged as a probe.
@@
From a model internals perspective, one can utilize techniques from interpretability research to find traces of local vs. global processing. For example, activation probing could be used to see if certain neurons or attention heads activate strongly when translating content that the model has seen before. If we had an LLM with known training data, we could take a sentence that appeared in training with its translation, present the source now, and compare the network activations to a case where we present a semantically similar sentence that never appeared in the “bilingual needles” during training. Do different heads light up? Does the model attend to different parts of its context or use different layers more intensively? Tools like integrated gradients or activation patching [^11] Activation patching was discussed in Section 5.4 above; see also Heimersheim and Nanda 2024. On a recent use of integrated gradients to analyze human and machine translation outputs, see Sizov et al. 2024. might let us “swap in” or suppress certain computations to test their effect on translation output. For instance, one could attempt to ablate the model’s high-level semantic representation (by intervening in middle layers) and see if it can still translate via literal mappings at lower layers, or vice versa. If disabling a semantic circuit causes the model to output word-by-word translations instead of fluent sentences, that would align with the two-system hypothesis (pointing to a distinct global semantic circuit). Conversely, knocking out a “memorization circuit” might degrade names and idiom translation while leaving general capability intact. These interpretability probes are speculative but increasingly feasible as researchers map circuits for specific abilities (for example, Wang et al. (2023) identified a circuit responsible for indirect object identification in GPT-2). A parallel effort could perhaps target a “bilingual lexicon circuit” versus a “semantic alignment circuit.”
@@ 7.3 Human Evaluation and Error Analysis
While automatic metrics will remain central, human evaluation can be indispensable in diagnosing how a translation was produced. Professional translators or bilingual speakers could examine model outputs for signs of literalness, creativity, and consistency with source. One experimental setup might involve a blind evaluation where annotators see pairs of translations from two versions of a model for the same input: a baseline model and one fine-tuned with selectively ablated data, as described in 7.1 above, or after priming the baseline model with parallel examples. If annotators consistently notice differences—say, one translation is more “word-for-word” and the other more “natural”—that would reveal the dual modes. One could also ask annotators to label errors in LLM translations by type: omission, addition, literal translation, mistranslation of idiom, terminology error, and other error categories. Alternatively, one could tailor the MQM framework to these purposes (Lommel et al., 2024). If the dual learning view is correct, we might find a bimodal error distribution: some errors look like over-literalness (the model sticking too closely to source form, potentially when it recalls a parallel), whereas others look like semantic drift or hallucination (the model trying to paraphrase meaning but going astray—an over-extension of global reasoning). Such patterns have been anecdotally reported by users and in case studies (e.g., models occasionally insert a plausible sentence that wasn’t in the source—a global inference gone wrong, or translate a proverb too directly—a “local memory” without adaptation). A systematic annotation of these phenomena across many outputs and languages would provide empirical grounding. Moreover, humans could probe the model through interactive evaluations: for instance, start translating a sentence and see if the model can continue (testing if it aligns with a human translator’s partial translation), or ask the model to explain its translation choices. If the model cites a source phrase mapping (e.g. “I translated X as Y because I’ve seen ‘X means Y’ before”) versus a reasoning (“X is similar to an English concept Z, so I rendered it as…”), it might indirectly reveal the influence of memorized pair vs. real-time reasoning. Of course, models can fabricate explanations, but with careful prompt engineering (or using decoder trace techniques), one might get useful signals.
@@ 7.4 Feasible Settings and Metrics
It’s worth considering what resources and scales are required for these investigations. Full retraining of a GPT-scale model for ablation is of course out of question, but fine-tuning experiments on smaller scale models (up to 7B or 13B parameters) are within reach for academic researchers today, especially with techniques like low-cost partial training (LoRA adapters or continual pre-training on smaller corpora). Using open-source LLMs like Llama or Mistral as bases would allow extraction of their training data sources to identify inherent parallel content, which is important for designing ablation studies.
Relatedly, this approarch could cast additional light on data contamination. For instance, one could compile all known parallel sentences in the RedPajama or Pile datasets and determine what portion of FLORES-200 test sets (Team et al., 2022) might have leaked; this is similar in spirit to the controlled study of contamination by Kocyigit et al. (2025), who examined how LLM evaluation on MT tasks is impacted when test examples overlap with training data. Leveraging such findings, we can intentionally create “contaminated” vs “clean” training splits and see how translation performance differs, thereby quantifying memorization.
In terms of evaluation metrics, beyond BLEU and chrF, metrics like COMET (which uses neural networks to predict quality) could be insightful since they might better capture adequacy vs. fluency differences. For measuring alignment resulting from Global learning, one might use cross-lingual retrieval tasks: e.g., compute the similarity of the model’s representations (from a certain layer) for a sentence and its translation. A model strong in global alignment should yield closer embeddings for translation pairs than for unrelated pairs. This could be evaluated by retrieval accuracy or clustering purity (an approach used, e.g., by Wang et al. 2022). If an ablated model (with no parallel fine-tuning data) still clusters translations together in embedding space, that’s evidence of global semantic alignment surviving without local signals. If a control model with added parallel data shows significantly tighter clustering, that quantifies the contribution of local learning. One might also repurpose the BLEU difference by shard metrics: divide a test set into segments where we believe the model had or lacked direct training exposure, and measure quality on each. A large gap would confirm the two modes.
@@ 8 Reconceptualizing Translation in the Age of Deep Learning
The possibility that high-performing translation by LLMs arises from dual origins prompts us to rethink what “translation” even means in this context. Historically, translation has never been a monolithic, one-size-fits-all process, neither for humans nor machines; and the rise of deep learning amplifies this pluralism. In this section, I consider some theoretical and practical implications of a dual-origin translation competence, touching on issues of pluralism in translation theory, emergent linguistic abilities, and the opacity of AI systems. I also draw connections to major shifts in translation studies and to the framework of distributional semantics that underlies modern NLP.
@@ 8.1 Pluralism and Duality in Translation
If LLMs indeed translate via two pathways—local recall vs. global inference, which are differentially informed by the corresponding learning processes—that underscores a broader lesson: there may be no singular Translation with a capital ‘T’. This resonates with the ethos of contemporary translation studies which emphasizes pluralism, the idea that what counts as translation can vary by its context, purpose, and agent (Snell-Hornby, 2006; Munday et al., 2022). Human translators adopt different strategies on the fly: sometimes translating word-for-word (e.g. for a legal text where fidelity to terminology is key), other times paraphrasing or explicating (e.g. when localizing a joke, an idiom, or a culturally-specific content for a new audience). The dualism proposal for LLMs suggests a similar internal diversity. Rather than a single unified algorithm, the model might be juggling something akin to the age-old dichotomy of “literal vs. free translation,” but doing so at a sub-symbolic level through patterns in data. From a theoretical perspective, this aligns with pluralist views that there is no one correct way to translate; success can be achieved through multiple different approaches, even within one mind (or one model).
@@ 8.2 Emergent Translation Competence?
One striking aspect of LLM translation is its emergent nature. The model’s ability to translate was not pre-programmed; it arose as a side effect of scale and data. This raises a question: to what extent is translation an emergent property of any sufficiently sophisticated information-processing system? If cross-lingual alignment can be achieved by exposure to multilingual data, then perhaps translation competence in general exists implicitly in the fabric of linguistic meaning. Philosophically, this touches on the concept of a universal meaning space, an idea with a long lineage (from Leinbiz’s characteristica universalis to the interlingua in MT research; see e.g. Hutchins 1986, Ch. 2). Modern distributional semantics gives this idea computational form: LLMs encode knowledge in high-dimensional vectors, and when trained on many languages, those vectors naturally form clusters by meaning rather than by language (at least for high-level abstractions). Empirical studies back this up: as mentioned earlier, mBERT, for example, was found to have aligned semantic representations across languages without any parallel data, a phenomenon noted as an “emerging cross-lingual structure” (Conneau et al., 2020b). Similarly, recent analyses show that certain LLM families have embeddings where tokens from different languages with the same meaning end up nearby, whereas other families’ embeddings cluster by language (Brinkmann et al., 2025; Lindsey et al., 2025). The fact that such structure can materialize in an artificial neural network suggests that translation is, at least partly, an emergent property of recognizing sameness of meaning across linguistic form.
@@
However, emergence has degrees. The dualism hypothesis suggests that what has emerged in LLMs is twofold: a capacity to detect translation equivalents when they are explicitly present (a kind of emergent memory of translation pairs), and a capacity to merge knowledge from different languages into one representation (an emergent “interlingua”). This dual emergence complicates our conceptualization. It means an LLM does not always deploy a single clean language-neutral representation; sometimes it relies on more language-specific correlations. In theoretical terms, we might say LLMs challenge the binary of “literal translation vs. interpretation” by doing both, depending on context. And they do so without specific instructions, unlike human translators who consciously choose strategies. This raises the following question: are we witnessing a rudimentary form of meta-cognition in the model (implicitly “deciding” between translation strategies based on input)? Or is it simply the probabilistic application of patterns? Either way, the outcome is an agent that can flexibly translate in different modes. In human terms, that would be considered a highly competent translator. So, should we credit the model with a form of translation competence? And if so, what does that competence consist in? Perhaps we need to take a cue from translation studies and broaden the definition of competence beyond “knows how to replace words in one language with words in another” (“translation as retrieval”) to “has internalized a cross-linguistic web of meaning that can be traversed in multiple ways” (“translation as inference”). This perspective connects with work in mainstream semantics suggesting that meaning is not tied to any single language, an idea that translation studies pioneers like Nida (1964) and linguists such as Jakobson (1959) would agree with in principle, though they never imagined it instantiated in billions of weighted connections.

### v3

@@ 7.1 Selective Ablation of Fine-Tuning Data
One direct way to probe the origins of LLMs’ translation behavior is to selectively remove or add certain data from an LLM’s training set and examine the impact. As outlined above (Section 6.2), this “retrospective surgery” approach was pioneered by briakou_searching_2023 on PaLM’s smaller replicas in a full-scale training regime. Could it be done in a less expensive fine-tuning mode and applied to other, increasingly available open-source models?
For example, one could take an existing pre-trained model, choose a new language pair, and continue training or fine-tuning the model on an additional dataset from which all parallel sentences for a new language pair are ablated, and then do the same with a version of that dataset where those pairs are exclusively retained but all the monolingual portions in the two chosen languages are removed. By comparing these two extremes—“no parallel data” vs. “only parallel data”—one could, in principle, observe differences in resulting translation performance. The dualism hypothesis predicts that a model fine-tuned without any parallel examples would still develop some translation ability in a new language pair (via global semantic alignment), but it might translate more slowly or less idiomatically for complex sentences. Conversely, a model fine-tuned only on parallel sentences would certainly learn direct mappings but might struggle on inputs that require world knowledge or cross-sentence context (since the training signal was confined to one-to-one sentence translations). If adding parallel data for, say, Russian-Japanese suddenly causes the model to produce very literal, consistent Japanese translations for Russian inputs (while translation outputs for other relevant language pairs remain unchanged), that would be a strong indication of a local learning effect.
Such controlled fine-tuning experiments could be done on smaller proxy models first, and data could be added incrementally, in both regimes. This might be similar to classic studies in cross-lingual word embeddings where parallel data were incrementally added to chiefly monolingual training to observe alignment effects (e.g. refining mBERT’s semantic space; cf. conneau_emerging_2020). Key data sources for real-world ablation studies include multilingual web crawls like mC4 or Common Crawl derivatives, where one can identify and remove bilingual segments (using language detectors and alignment algorithms). However, smaller corpus resources—from the latest MT benchmarks or recent real-life translation projects (e.g., the Reeve Corpus, balashov-etal-2025-translation)—could also be used in the fine-tuning mode. Model sizes should vary: experiments on 1B, 5B, and 10B parameter models can indicate scaling trends.
@@
A note of caution is warranted here. The preceding considerations regarding controlled ablation strictly apply to full-scale fine-tuning, in which all model parameters are updated under the same regime as in the initial training. This process is computationally expensive, even for relatively small (1–10B parameter) open-source LLMs. In practice, highly efficient Low-Rank Adaptation (LoRA) methods have become standard for fine-tuning across many tasks, including translation. [^11] See Sections 3.5, 4.2, 5.1, and references therein. While the practical value of LoRA and related techniques is indisputable, their use in testing the duality hypothesis outlined above may be problematic. Our objective is to assess the differential impact of two types of fine-tuning data—those with and without “bilingual needles”—on the emergence of highly specific translation abilities in fine-tuned LLMs. With LoRA, however, the fine-tuning signal is diverted away from the main model weights (which remain frozen) into a set of additional low-rank matrices. This architectural change precludes informed judgments about the comparative roles of Local and Global learning, as all new learning is now channeled into the adaptation weights. This redirection may produce a “third pathway” for developing specialized translation abilities, one that merits investigation in its own right; but its relevance to the Local-Global learning dilemma remains uncertain. In light of these considerations, full-scale fine-tuning, though more resource-intensive, appears to be the more appropriate approach.
@@ 7.2 Synthetic Tasks and Controlled Probes
Another strategy is to design diagnostic tasks that disentangle local from global translation capabilities without the need for retraining or even fine-tuning. An intriguing direction is to investigate whether the distinction between local and global mechanisms re-emerges within the framework of in-context learning.
One way to approach this question is to provide the model with synthetic bilingual contexts at inference time and examine how it exploits these cues. For instance, we could feed the model a prompt that includes a few embedded parallel examples for a new language pair (e.g. a mini-glossary or several translated sentences) and then ask it to translate a novel sentence. This is akin to few-shot prompting. If local learning is a distinct operative mechanism, the model should improve or change style given those explicit examples, essentially performing cross-lingual “translation memory” look-up. On the other hand, if it relies mostly on global knowledge, a few extra examples might not change its performance much. By varying the content of these examples (literal translations vs. paraphrastic ones) and observing the output, one can infer the model’s default bias. A model heavily skewed to local learning might even mistranslate by over-following the format of a misleading example (e.g., if given a bad literal translation as a cue, it might mimic that literality). In contrast, a globally oriented model would stick closer to conveying meaning and be less perturbed by small prompt changes. Recent work on prompt-based analysis of LLM translation indicates that certain prompt formats act as “soft switches” for translation style (aldosari_assessing_2025); this can be leveraged as a probe.
@@
From a model internals perspective, one can utilize techniques from interpretability research to find traces of local vs. global processing. For example, activation probing could be used to see if certain neurons or attention heads activate strongly when translating content that the model has seen before. If we had an LLM with known training data, we could take a sentence that appeared in training with its translation, present the source now, and compare the network activations to a case where we present a semantically similar sentence that never appeared in the “bilingual needles” during training. Do different heads light up? Does the model attend to different parts of its context or use different layers more intensively (zhang_exploring_2025)? Tools like integrated gradients or activation patching [^12] Activation patching was discussed in Section 5.4 above; see also heimersheim_how_2024. On a recent use of integrated gradients to analyze human and machine translation outputs, see sizov-etal-2024-analysing. might let us “swap in” or suppress certain computations to test their effect on translation output. For instance, one could attempt to ablate the model’s high-level semantic representation (by intervening in middle layers) and see if it can still translate via literal mappings at lower layers, or vice versa. If disabling a semantic circuit causes the model to output word-by-word translations instead of fluent sentences, that would align with the duality hypothesis (pointing to a distinct global semantic circuit). Conversely, knocking out a “memorization circuit” might degrade names and idiom translation while leaving general capability intact. These interpretability probes are speculative but increasingly feasible as researchers map circuits for specific abilities. [^13] For example, wang_interpretability_2023 earlier identified a circuit responsible for indirect object identification in GPT-2. A parallel effort could perhaps target a “bilingual lexicon circuit” versus a “semantic alignment circuit.”
@@ 7.3 Human Evaluation and Error Analysis
While automatic metrics will remain central, human evaluation can be indispensable in diagnosing how a translation was produced. Professional translators or bilingual speakers could examine model outputs for signs of literalness, creativity, and consistency with source. One experimental setup might involve a blind evaluation where annotators see pairs of translations from two versions of a model for the same input: a baseline model and one fine-tuned with selectively ablated data, as described in Section 7.1 above, or after priming the baseline model with parallel examples (Section 7.2). If annotators consistently notice differences—say, one translation is more “word-for-word” and the other more “natural”—that would reveal the dual modes. One could also ask annotators to label errors in LLM translations by type: omission, addition, literal translation, mistranslation of idiom, terminology error, and other error categories. Alternatively, one could tailor the MQM framework to these purposes (lommel-etal-2024-multi) and experiment with different context spans—the number of preceding and/or succeeeding sentences shown to the annotators (castilho_survey_2025). If the dual learning view is correct, we might find a bimodal error distribution: some errors look like over-literalness (the model sticking too closely to source form, potentially when it recalls a parallel), whereas others look like semantic drift or hallucination (the model trying to paraphrase meaning but going astray—an over-extension of global reasoning). Such patterns have been anecdotally reported by users and in case studies (e.g., models occasionally insert a plausible sentence that wasn’t in the source—a global inference gone wrong, or translate a proverb too directly—a “local memory” without adaptation). A systematic annotation of these phenomena across many outputs and languages would provide empirical grounding. Moreover, human experts could probe the model through interactive evaluations: for instance, start translating a sentence and see if the model can continue (testing if it aligns with a human translator’s partial output), or ask the model to explain its translation choices. If the model cites a source phrase mapping (e.g. “I translated X as Y because I’ve seen ‘X means Y’ before”) versus a reasoning (“X is similar to an English concept Z, so I rendered it as…”), it might indirectly reveal the influence of memorized pair vs. real-time reasoning. Of course, models can fabricate explanations, but with careful prompt engineering (or using decoder trace techniques), one might get useful signals.
@@ 7.4 Feasible Settings and Metrics
It’s worth considering what resources and scales are required for these investigations. Full retraining of a GPT-scale model for ablation is of course out of question, but fine-tuning experiments on smaller scale models (up to 7B or 13B parameters) are within reach for academic researchers today. Using open-source LLMs like Llama, Mistral, OLMo (olmo20252olmo2furious), or GPT-OSS (openai_introducing_2025) as bases would allow extraction of their training data sources to identify inherent parallel content, which is important for designing ablation studies.
Relatedly, this approarch could cast additional light on data contamination. For instance, one could compile all known parallel sentences in the RedPajama or Pile datasets and determine what portion of FLORES-200 test sets (team_no_2022) might have leaked; this is similar in spirit to the controlled study of contamination by kocyigit_overestimation_2025, who examined how LLM evaluation on MT tasks is impacted when test examples overlap with training data. Leveraging such findings, we can intentionally create “contaminated” vs “clean” training splits and see how translation performance differs, thereby quantifying memorization.
For automatic evaluation, string-based metrics (e.g., BLEU, chrF) can flag potential “memorized translations,” whereas neural metrics such as COMET may be more informative because they better capture differences between adequacy and fluency. For measuring alignment resulting from Global learning, one might use cross-lingual retrieval tasks: e.g., compute the similarity of the model’s representations (from a certain layer) for a sentence and its translation. A model strong in global alignment should yield closer embeddings for translation pairs than for unrelated pairs. This could be evaluated by retrieval accuracy or clustering purity. [^14] An approach used, e.g., by wang_english_2022. Cf. earlier work on reference-free neural network-based evaluation of monolingual inter-sentential “cohesiveness” showing the effectiveness of even very simple earlier static embeddings such as Work2Vec (elmakias_oblivious_2021). If an ablated model (with no parallel fine-tuning data) still clusters translations together in embedding space, that’s evidence of global semantic alignment surviving without local signals. If a control model with added parallel data shows significantly tighter clustering, that quantifies the contribution of local learning. One might also repurpose the BLEU difference by shard metrics: divide a test set into segments where we believe the model had or lacked direct training exposure, and measure quality on each. A large gap would confirm the two modes.
@@ 8 Reconceptualizing Translation in the Age of Deep Learning
The possibility that high-performing translation by LLMs arises from dual origins prompts us to rethink what “translation” even means in this context. Historically, translation has never been a monolithic, one-size-fits-all process, neither for humans nor machines; and the rise of deep learning amplifies this pluralism. In this section, I consider broader theoretical implications of a dual-origin translation competence, touching on issues of pluralism in translation theory, emergent linguistic abilities, and the opacity of AI systems. I also draw connections to major shifts in translation studies, to the undelying ideas in classical philosophy of translation, and to the framework of distributional semantics that underlies modern NLP.
@@ 8.1 Pluralism and Duality in Translation
If LLMs indeed translate via two pathways—local recall vs. global inference, which are differentially informed by the corresponding learning processes—that underscores a broader lesson: there may be no singular Translation with a capital ‘T’. This resonates with the ethos of contemporary translation studies which emphasizes pluralism, the idea that what counts as translation can vary by its context, purpose, and agent (snell-hornby_turns_2006; munday_introducing_2022). Human translators adopt different strategies on the fly: sometimes translating word-for-word (e.g. for legal texts or patent materials where fidelity to terminology is key), other times paraphrasing or explicating (e.g. when localizing a joke, an idiom, or a culturally-specific content for a new audience). The dualism proposal for LLMs suggests a similar internal diversity. Rather than a single unified algorithm, the model might be juggling something akin to the age-old dichotomy of “literal vs. free translation,” but doing so at a sub-symbolic level through patterns in data. From a theoretical perspective, this aligns with pluralist views that there is no one correct way to translate; success can be achieved through multiple different approaches, even within one mind (or one model).
@@
This pluralism also resonates with classic philosophical treatments of translation. On Quine’s view, multiple “translation manuals” can fit all the behavioral (and textual) evidence equally well; indeterminacy of translation is not failure but a structural feature of the process (quine_word_1960). Davidson likewise frames translation as radical interpretation: recovering meaning by imposing rational coherence (his Principle of Charity) on a speaker’s utterances in context (davidson_radical_1973). Read through this lens, the local/global duality in LLMs is a computational echo of those frameworks: “local recall” resembles a manual-like mapping, while “global alignment” resembles the Davidsonian drive toward holistic coherence across a broad web of usage. Both modes can yield adequate translations, yet neither is uniquely mandatory. One can think of it as a modern, sub-symbolic instance of pluralism anticipated in classical philosophy of translation. [^15] See rawling_routledge_2019 for a recent collection of articles exploring important questions that arise at the interface of classical translation studies and philosophy of language.
@@ 8.2 Emergent Translation Competence?
One striking aspect of LLM translation is its emergent nature. The model’s ability to translate was not pre-programmed; it arose as a side effect of scale and data. This raises a question: to what extent is translation an emergent property of any sufficiently sophisticated information-processing system? If cross-lingual alignment can be achieved by exposure to multilingual data, then perhaps translation competence in general exists implicitly in the fabric of linguistic meaning. Philosophically, this touches on the concept of a universal meaning space, an idea with a long lineage (from Leinbiz’s characteristica universalis to the interlingua in MT research; see e.g. hutchins_machine_1986, Ch. 2). Modern distributional semantics gives this idea computational form: LLMs encode knowledge in high-dimensional vectors, and when trained on many languages, those vectors naturally form clusters by meaning rather than by language (at least for high-level abstractions). Empirical studies back this up: as mentioned earlier, mBERT, for example, was found to have aligned semantic representations across languages without any parallel data, a phenomenon noted as an “emerging cross-lingual structure” (conneau_emerging_2020). Similarly, recent analyses show that certain LLM families have embeddings where tokens from different languages with the same meaning end up nearby, whereas other families’ embeddings cluster by language (brinkmann-etal-2025-large; lindsey_biology_2025). The fact that such structure can materialize in an artificial neural network suggests that translation is, at least partly, an emergent property of recognizing sameness of meaning across linguistic form.
quine_word_1960 cautions us against reifying this emergent common space as a uniquely correct “interlingua.” If translation manuals are underdetermined by evidence, then a high-dimensional distributional space may encode many near-equivalent alignments compatible with the same corpus. LLMs select among them based on their training mix and objective. A Davidsonian moral (davidson_radical_1973) might be somewhat different: successful interpretation is secured by maximizing truth and coherence across a speaker’s total (or model-internal) corpus. In practice, instruction-following and decoding objectives act as “charity-like” pressures, steering outputs toward globally coherent continuations. Thus, along one dimension of this classical philosophical framework (Quine), we should expect principled underdetermination within the model’s latent alignments; along another (Davidson), we could expect convergence in use, because the system is constantly optimized to produce coherent, truth-cunducive continuations under task constraints.
@@
However, emergence has degrees. The dualism hypothesis suggests that what has emerged in LLMs is twofold: a capacity to detect translation equivalents when they are explicitly present (a kind of emergent memory of translation pairs), and a capacity to merge knowledge from different languages into one representation (an emergent “interlingua”). This dual emergence complicates our conceptualization. It means an LLM does not always deploy a single clean language-neutral representation; sometimes it relies on more language-specific correlations. In theoretical terms, we might say LLMs challenge the binary of literal translation vs. “transcreation” by doing both, depending on context. And they do so without specific instructions, unlike human translators who consciously choose strategies. This raises the following question: are we witnessing a rudimentary form of meta-cognition in the model (implicitly “deciding” between translation strategies based on input)? Or is it simply the probabilistic application of patterns? Either way, the outcome is an agent that can flexibly translate in different modes. In human terms, that would be considered a highly competent translator. So, should we credit the model with a form of translation competence? And if so, what does that competence consist in? Perhaps we need to take a cue from translation studies and broaden the definition of competence beyond “knows how to replace words in one language with words in another” (“translation as retrieval”) to “has internalized a cross-linguistic web of meaning that can be traversed in multiple ways” (“translation as inference”). This perspective connects with work in mainstream semantics suggesting that meaning is not tied to any single language, an idea that translation studies pioneers like nida_toward_1964 and linguists such as jakobson_linguistic_1959 would agree with in principle, though they never imagined it instantiated in billions of weighted connections.